# Privacy-Preserving Retrieval Augmented Generation with Differential Privacy

### v1

@@ 3.2 DPSparseVoteRAG: Differentially Private Voting Algorithm for RAG with Sparse Vector Technique
Input: Prompt $x$, External data source $D$, Generator $\mathrm{LLM}$, Retriever $R$, # of voters $m$, # of retrieval per voter $k$, Per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$, Total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$, Threshold $\tau$
Output: Private answer $y$
1
@@
3 $D_{1},\ldots,D_{m}\leftarrow$ split $D$ into $m$ disjoint subset;
4
5$(\epsilon_{\mathrm{token-RAG}},\delta_{\mathrm{token-RAG}})\leftarrow$ $(\epsilon_{\mathrm{token}}/2,\delta_{\mathrm{token}})$;
6 $\epsilon_{\mathrm{token-Lap}}\leftarrow\epsilon_{\mathrm{token}}/2$;
7 $c\leftarrow 1$;
8 $t\leftarrow 1$;
9 $\tau_{1}\leftarrow\tau+\mathrm{Lap}(\nicefrac{{2}}{{\epsilon_{\mathrm{token-Lap}}}})$;
10 while $c\leq c_{\max}$ do
11       $y_{t}^{\mathrm{non-RAG}}\leftarrow\mathrm{LLM}_{t}(x,``^{\prime\prime},y_{<t})$;
12       for $i\leftarrow 1$ to $m$ do
             $y_{t}^{(i)}\leftarrow\mathrm{LLM}_{t}(x,R(x,D_{i}),y_{<t})$;
              // Generate a token with RAG
13
14       end for
15      $h_{t}\leftarrow$ Build a histogram of tokens from $y_{t}^{(1)},\ldots,y_{t}^{(m)}$;
16       $a_{t}\leftarrow$ Extract a count of $h_{t}$ at $y_{t}^{\mathrm{non-RAG}}$;
17       $\nu_{t}\leftarrow\mathrm{Lap}(\nicefrac{{4}}{{\epsilon_{\mathrm{token-Lap}}}})$;
18       if $a_{t}+\nu_{t}\leq\tau_{c}$ then
             /* Privately choose the most frequent token when RAG outputs disagree with non-RAG output */
19             $y_{t}\leftarrow\mathrm{LimitedDomain}(h_{t},\epsilon_{\mathrm{token-RAG}},\delta_{\mathrm{token-RAG}})$;
20             $c\leftarrow c+1$;
21             $\tau_{c}\leftarrow\tau+\mathrm{Lap}(\nicefrac{{2}}{{\epsilon_{\mathrm{token-Lap}}}})$;
22
23       end if
24      else
             /* Use non-RAG output when RAG outputs agree with it */
25             $y_{t}\leftarrow y_{t}^{\mathrm{non-RAG}}$;
26
27       end if
[…]
             return $(y_{1},\ldots,y_{t})$;
              // Halt if end of sequence
29
30       end if
31      $t\leftarrow t+1$;
32
33 end while
34return $(y_{1},\ldots,y_{t})$;
Algorithm 2 DPSparseVoteRAG
The main drawback of the aforementioned algorithm is that we need to spend a non-negligible amount of privacy budget for each token to guarantee its quality. This prevents our algorithm to generate longer answers—sometimes it can halt before it generates the actual answers due to privacy budget shortage. More concretely, consider the following question-answering example:
@@ 31      $t\leftarrow t+1$;
However, having a closer look to our voting algorithm, we observe that there is a room for improvement. When generating the 3rd word, Gatsby, every input of the LLM contains The Great, 1st and 2nd previously output words, and the great gatsby, a part of the question, even though the provided retrieved documents are different. Thus, the LLM should successfully generate Gatsby without access to the sensitive information. Ideally, we should not spend a privacy budget for such a word.
We address this by incorporating the sparse vector technique into our voting algorithm, yielding our improved algorithm DPSparseVoteRAG. In particular, before we apply private voting among generated tokens, we check if the generated tokens coincide with the token generated by the LLM without retrieved documents appended, i.e., the input is composed of the prompt and previously generated tokens only. We continue to the voting only when they do not coincide. Otherwise, we use the LLM output without retrieved documents. It is shown from the analysis of the sparse vector technique that the consumed privacy budget scales with the number of times that it uses the private voting, not with the total number of generated tokens. Consequently, the resulting algorithm, shown in Algorithm 2, enables us to spend privacy budget only when it needs sensitive information. We note that as a result of this change in our algorithm, we compute the maximum number of tokens to generate with private voting, $c_{\max}$, from the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$, instead of the maximum number of token generation, $T_{\max}$, as in DPVoteRAG but in the same way to compute $T_{\max}$. The formal privacy analysis of this algorithm is as follows. The guarantee holds due to the privacy guarantee of the LimitedDomain mechanism and the NumericSparse algorithm [12].
@@ 4.1 Methodology
We use two question-answering benchmarking datasets for RAG: Trivia [18] and Natural Question (NQ) [21]. Each dataset consists of a list of pairs of question and answer lists, i.e., every question can have multiple answers. By following the standard evaluations in RAG [5, 23, 16], we use the Wikipedia dataset as the external data source from which a retriever finds relevant documents. For each dataset, we use a subset of 100 questions to manage the computational overhead  [^4] We first filter out questions with less than 20 ground truth documents in the Wikipedia dataset. If a question relates with only a few documents, it is inevitable that DP algorithms fail since replacing one document would change the output a lot. Then, we split remaining questions into 4 bins with 20–29, 30–39, 40–49, and 50–59 relevant documents and sample 25 questions from each bin..
@@
The retriever we use is the Dense Passage Retriever (DPR) [20] which is built on top of BERT [6]. The DPR finds relevant documents that are close to the question query in the embedding space produced by BERT. We compare the following generator LLMs: OPT (1.3B) [41], GPT2-XL [31], and Pythia (1.4B) [3]. We additionally report the result of OPT (2.7B) in Appendix A.1.
@@
We observe the results under multiple total privacy budgets, $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. More specifically, we sweep $\epsilon_{\mathrm{total}}=2$ to $40$ and set $\delta_{\mathrm{total}}=10^{-4}$. Furthermore, we consider different per-token privacy budgets for our private algorithms: $\epsilon_{\mathrm{token}}=1,2,5$ and $\delta_{\mathrm{token}}=10^{-5}$. We consider the number of voters $m$ of $10$, $20$, $30$, $40$, and $50$ for VoteRAG, and $30$, $40$, and $50$ for DPVoteRAG and DPSparseVoteRAG so as to ensure reasonable privacy-utility tradeoff and computational overhead. Too small number of voters yields too large $\epsilon_{\mathrm{token}}$ to generate accurate answers. Increasing the number of voters not only increases computational overhead but also requires a lot of relevant documents which eventually include non-relevant ones, resulting in poor utility. For DPSparseVoteRAG, we set the threshold $\tau$ to be the half of the number of voters, i.e, $\tau=m/2$. When we use the LimitedDomain mechanism to privately choose the most frequent token, we set their parameter $\bar{k}$ to be the number of voter, where $\bar{k}$ is the limited size of the domain to which we add the Gumbel noise. For voting algorithms, each voter receives 1 relevant document from corresponding subset of external knowledge dataset $D$. The utility evaluation metric is the match accuracy [25, 2, 33, 42] which measures if the prediction to a question contains any of its answers.
@@ 4.2 Results
Figure 3: Average match accuracy comparison across algorithms on Trivia (upper row) and NQ (lower row) datasets with different generator LLMs: OPT (1.3B) (left column), GPT2-XL (middle column), and Pythia (1.4B) (right column). We report the best results over hyperparameters for each $\epsilon_{\mathrm{total}}$. Note that the performances of Non-RAG and VoteRAG on Trivia with Pythia (1.4B) coincide (upper right).
Comparison of algorithms.
Figure 3 shows the average match accuracy of baseline algorithms, Non-RAG and Vote-RAG, and our algorithms, DPVoteRAG and DPSparseVoteRAG for different total privacy guarantees ($\epsilon_{\mathrm{total}}$). We first observe that VoteRAG mostly improves the utility compared to Non-RAG suggesting the benefit of RAG for knowledge-extensive question-answering tasks. The only exception is the one with Pythia on Trivia—VoteRAG and Non-RAG perform comparably. We anticipate this is because Pythia is trained on Wikipedia passages and memorizes them well. Furthermore, across different datasets and generator LLMs, we observe that DPSparseVoteRAG outperforms Non-RAG mostly under $\epsilon_{\mathrm{total}}\geq 10$ when VoteRAG is better than Non-RAG. On the other hand, DPVoteRAG requires larger total privacy budget to outperform Non-RAG or can even perform worse than Non-RAG. Comparing the utility between our algorithms, we see DPSparseVoteRAG always yields better results than DPVoteRAG. This suggests the importance of the smart privacy budget usage offered by DPSparseVoteRAG.
Table 1: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$.
@@
$\epsilon_{\mathrm{token}}=1$ 0.33 0.37 0.38 0.39 0.45 0.45 0.39 0.45 0.45 0.39 0.45 0.45
$\epsilon_{\mathrm{token}}=2$ 0.06 0.05 0.05 0.41 0.42 0.40 0.49 0.49 0.47 0.49 0.49 0.47
$\epsilon_{\mathrm{token}}=5$ 0.00 0.00 0.00 0.06 0.05 0.05 0.34 0.35 0.34 0.47 0.47 0.46
@@
$\epsilon_{\mathrm{token}}=1$ 0.30 0.41 0.40 0.30 0.41 0.40 0.30 0.41 0.40 0.30 0.41 0.40
$\epsilon_{\mathrm{token}}=2$ 0.38 0.43 0.45 0.45 0.51 0.55 0.46 0.54 0.58 0.46 0.54 0.58
$\epsilon_{\mathrm{token}}=5$ 0.22 0.23 0.26 0.47 0.48 0.47 0.53 0.54 0.52 0.60 0.59 0.58
Effects of hyperparameters.
We next investigate the effects of hyperparameters of our algorithms by taking a closer look at the results in Table 1 with OPT (1.3B) on Trivia dataset under different total privacy budgets $\epsilon_{\mathrm{total}}$. We provide the detailed results with other LLMs on both Trivia and NQ in Appendix A.1.
Commonly between DPVoteRAG and DPSparseVoteRAG, we observe that the optimal $\epsilon_{\mathrm{token}}$ increases as we allow more total privacy budgets. For a tight total privacy budget, large $\epsilon_{\mathrm{token}}$ allows our algorithms to only output a few meaningful tokens; thus, smaller $\epsilon_{\mathrm{token}}$ is preferable. Conversely, under a large total privacy budget, accurate token generation with large $\epsilon_{\mathrm{token}}$ is more important than having more tokens generated with small $\epsilon_{\mathrm{token}}$. The main difference between two of our algorithms in terms of $\epsilon_{\mathrm{token}}$ is that DPSparseVoteRAG generally allows us to set larger $\epsilon_{\mathrm{token}}$ than DPVoteRAG under a fixed total privacy budget. More specifically, while setting $\epsilon_{\mathrm{token}}=2$ yields the best utility for DPVoteRAG even under $\epsilon_{\mathrm{total}}=40$, DPSparseVoteRAG performs the best with $\epsilon_{\mathrm{token}}=5$. This demonstrates the benefit of DPSparseVoteRAG to save and spend privacy budget cleverly—it can spend the saved privacy budget for generating important tokens for answering questions correctly.
With regard to the number of voters $m$ for DPVoteRAG, we generally see that having more voters results in better utility with $\epsilon_{\mathrm{token}}=1$. On the other hand, with larger $\epsilon_{\mathrm{token}}$, the number of voters has less effect on the utility and having less voters provides slightly better utility. The observation holds for all values of $\epsilon_{\mathrm{total}}$. This is due to the two distinct consequences of having more voters. More voters alleviate the effect of DP noise to the token histograms constructed within the algorithm. However, depending on the number of relevant documents to the question, there is a risk of having voters with irrelevant documents who can vote for wrong tokens. The first consequence is more dominant particularly under small $\epsilon_{\mathrm{token}}$ while the second one becomes more dominant under larger $\epsilon_{\mathrm{token}}$. For DPSparseVoteRAG, although we observe the similar trend to the above, having more voters is effective until $\epsilon_{\mathrm{token}}=2$ as opposed to $\epsilon_{\mathrm{token}}=1$ for DPVoteRAG. This is because DPSparseVoteRAG uses one half of $\epsilon_{\mathrm{token}}$ for the token generation when the RAG outputs disagree with the non-RAG output—we see the consistent impacts of the number of voters across algorithms.
4.3 Discussion
@@
Our thorough empirical evaluation has revealed that our algorithms improves the utility of question-answering tasks over the non-RAG baseline by exploiting the external knowledge through RAG while ensuring a reasonable level of privacy $\epsilon_{\mathrm{total}}=10$. While this is mostly the case, we also observe that even (non-private) VoteRAG does not improve the utility on Trivia with Pythia. This observation implies that RAG and its private versions need to use truly sensitive data, which is completely outside the pre-training and fine-tuning data of LLMs, to be valuable enough. On the flip side, we anticipate our algorithms are even more effective for more involved tasks which heavily require the external data source for LLMs to generate meaningful responses as it is expected non-RAG LLMs perform poorly for such tasks.
@@
We further find that DPSparseVoteRAG is consistently outperforming DPVoteRAG across different LLMs and datasets. DPSparseVoteRAG augments DPVoteRAG by utilizing the non-RAG LLM and the sparse vector technique so that it only spends a privacy budget for an output token requiring sensitive external knowledge. The consistently better performances of DPSparseVoteRAG suggest the importance of separately treating token generations for meaningful tokens, i.e., tokens requiring external knowledge, and for other general tokens in the privacy-constraint setting.
@@
The detailed experimental results across different values of the hyperparameters give us some guidelines on how to set them properly. In particular, the per-token privacy budget, $\epsilon_{\mathrm{token}}$, should be small enough to allow the algorithms to generate a reasonably large number of tokens ($\sim 10$). As far as the resulting sequence is long enough, we should certainly make it as large as possible to enable accurate token generations. Furthermore, the number of voters, $m$, should be set by considering per-token generation quality. For a strict per-token privacy budget, e.g., $\epsilon_{\mathrm{token}}=1$, it needs to be large enough to relax the impact of the DP noise to the vote result histogram. For larger per-token privacy budgets, it should be small enough to ensure the majority of the voters receives useful external documents as the context.

### v2

@@ 3.2 DPSparseVoteRAG: Differentially Private Voting Algorithm for RAG with Sparse Vector Technique
Algorithm 2 DPSparseVoteRAG
0:  Prompt $x$, External data source $D$, Generator $\mathrm{LLM}$, Retriever $R$, # of voters $m$, # of retrieval per voter $k$, Per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$, Total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$, Threshold $\tau$, Maximum # of output tokens (regardless of privacy) $T_{\max}$
0:  Private answer $y$
1:  // Privacy budget setup
2:  $(\epsilon_{\mathrm{token-RAG}},\delta_{\mathrm{token-RAG}})\leftarrow(\epsilon_{\mathrm{token}}/2,\delta_{\mathrm{token}})$, $\epsilon_{\mathrm{token-Lap}}\leftarrow\epsilon_{\mathrm{token}}/2$
3:  $c_{\max}\leftarrow$ maximum # of tokens to generate privately based on $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$
4:  $c\leftarrow c_{\max}$, $\hat{\tau}\leftarrow\tau+\mathrm{Lap}(\nicefrac{{2}}{{\epsilon_{\mathrm{token-Lap}}}})$
5:  // Retrieval and random partition of the relevant documents
6:  $D_{x}\leftarrow$ Retrieve $mk$ most relevant documents $R(x,D;mk)$.
7:  $D_{x}^{1},\cdots,D_{x}^{m}\leftarrow$ Uniformly randomly partition $D_{x}$ into $m$ disjoint subsets.
8:  for $t\leftarrow 1$ to $T_{\max}$ do
9:     // Generating the non-private token and token histogram with the sensitive documents
10:     $y_{t}^{\mathrm{non-RAG}}\leftarrow\mathrm{LLM}_{t}(x,``^{\prime\prime},y_{<t})$
11:     for $i\leftarrow 1$ to $m$ do
12:        $y_{t}^{(i)}\leftarrow\mathrm{LLM}_{t}(x,D_{x}^{i},y_{<t})$
13:     end for
14:     $h_{t}\leftarrow$ Build a histogram of tokens from $y_{t}^{(1)},\ldots,y_{t}^{(m)}$
15:     // Producing a token from the histogram privately only when $y_{t}^{\mathrm{non-RAG}}$ is uncommon in $h_{t}$
16:     $a_{t}\leftarrow$ Extract a count of $h_{t}$ at $y_{t}^{\mathrm{non-RAG}}$
17:     if $a_{t}+\mathrm{Lap}(\nicefrac{{4}}{{\epsilon_{\mathrm{token-Lap}}}})\leq\hat{\tau}$ then
18:        $y_{t}\leftarrow\mathrm{LimitedDomain}(h_{t},\epsilon_{\mathrm{token-RAG}},\delta_{\mathrm{token-RAG}})$
19:        // The privacy budget is only consumed when $y_{t}$ is from the histogram
20:        $c\leftarrow c-1$, $\hat{\tau}\leftarrow\tau+\mathrm{Lap}(\nicefrac{{2}}{{\epsilon_{\mathrm{token-Lap}}}})$
21:     else
22:        $y_{t}\leftarrow y_{t}^{\mathrm{non-RAG}}$
23:     end if
24:     // Halting if end of sequence or the privacy budget has been exhausted
@@
26:        return $(y_{1},\ldots,y_{t})$
27:     end if
28:  end for
29:  return $(y_{1},\ldots,y_{T_{\max}})$
The main drawback of the aforementioned algorithm is that we need to spend a non-negligible amount of privacy budget for each token to guarantee its quality. This prevents our algorithm from generating longer answers—sometimes it can halt before it generates the actual answers due to privacy budget shortage. More concretely, consider the following question-answering example:
@@
However, having a closer look at our voting algorithm, we observe that there is room for improvement. When generating the 3rd word, Gatsby, every input of the LLM contains The Great, 1st and 2nd previously output words, and the great gatsby, a part of the question, even though the provided retrieved documents are different. Thus, the LLM should successfully generate Gatsby without access to the sensitive information. Ideally, we should not spend a privacy budget for such a word.
We address this by incorporating the sparse vector technique into our voting algorithm, yielding our improved algorithm DPSparseVoteRAG. In particular, before we apply private voting among generated tokens, we check if the generated tokens coincide with the token generated by the LLM without retrieved documents appended, i.e., the input is composed of the prompt and previously generated tokens only. We continue to the voting only when they do not coincide. Otherwise, we use the LLM output without retrieved documents. It is shown from the analysis of the sparse vector technique that the consumed privacy budget scales with the number of times that it uses the private voting, not with the total number of generated tokens. Consequently, the resulting algorithm, shown in Algorithm 2, enables us to spend a privacy budget only when it needs sensitive information. We note that as a result of this change in our algorithm, we compute the maximum number of tokens to generate with private voting, $c_{\max}$, from the per-token privacy budget $(\epsilon_{\mathrm{token}},\delta_{\mathrm{token}})$ and total privacy budget $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$, instead of the maximum number of token generation, $T_{\max}$, as in DPVoteRAG but in the same way to compute $T_{\max}$. The formal privacy analysis of this algorithm is as follows. The guarantee holds due to the privacy guarantee of the LimitedDomain mechanism and the NumericSparse algorithm [13].
@@ 4.1 Methodology
We use two question-answering benchmarking datasets for RAG: Trivia [19] and Natural Question (NQ) [22]. Each dataset consists of a list of pairs of question and answer lists, i.e., every question can have multiple answers. By following the standard evaluations in RAG [5, 24, 17], we use the Wikipedia dataset as the external data source from which a retriever finds relevant documents. For each dataset, we use a subset of 100 questions to manage the computational overhead  [^4] We first filter out questions to less than 20 ground truth documents in the Wikipedia dataset. If a question relates with only a few documents, DP algorithms will likely fail since replacing a document would change the output a lot. Then, we split the remaining questions into 4 bins with 20–29, 30–39, 40–49 and 50–59 relevant documents and sample 25 questions from each bin..
@@
The retriever we use is the Dense Passage Retriever (DPR) [21] which is built on top of BERT [6]. It finds relevant documents that are close to the question in the embedding space produced by BERT. We compare the following generator LLMs: OPT (1.3B) [42], Llama 3.1 (8B) [8], and Pythia (1.4B) [3]. We additionally report the result of OPT (2.7B), Llama 3.2 (1B), and GPT2-XL [32] in Appendix B.1.
@@
We observe the results under multiple total privacy budgets, $(\epsilon_{\mathrm{total}},\delta_{\mathrm{total}})$. More specifically, we sweep $\epsilon_{\mathrm{total}}=2$ to $40$ and set $\delta_{\mathrm{total}}=10^{-4}$. Furthermore, we consider different per-token privacy budgets for our private algorithms: $\epsilon_{\mathrm{token}}=1,2,5$ and $\delta_{\mathrm{token}}=10^{-5}$. We consider the number of voters $m$ of $10$, $20$, $30$, $40$, and $50$ for VoteRAG, and $30$, $40$, and $50$ for DPVoteRAG and DPSparseVoteRAG so as to ensure reasonable privacy-utility tradeoff and computational overhead. For DPSparseVoteRAG, we set the threshold $\tau$ to be half of the number of voters, i.e., $\tau=m/2$. When we use the LimitedDomain mechanism to privately choose the most frequent token, we set their parameter $\bar{k}$ to be the number of voters, where $\bar{k}$ is the limited size of the domain to which we add the Gumbel noise. For voting algorithms, each voter receives 1 relevant document, i.e., $k=1$. The utility evaluation metric is the match accuracy [26, 2, 34, 43] which measures if the prediction to a question contains any of its answers.
@@ 4.2 Results
Figure 3: Average match accuracy comparison across algorithms on Trivia (upper row) and NQ (lower row) datasets with different generator LLMs: OPT (1.3B) (left column), Pythia (1.4B) (middle column), and Llama 3.1 (8B) (right column). The reported results are the means and standard deviations of average match accuracy over three runs. We report the best results over hyperparameters for each $\epsilon_{\mathrm{total}}$.
Table 1: Average match accuracy comparison of our algorithms on Trivia dataset with OPT (1.3B) under varying values of total privacy budget $\epsilon_{\mathrm{total}}$ with different hyperparameters, the number of voters $m$ and $\epsilon_{\mathrm{token}}$. Bold font values represent the best performance of each algorithm under fixed $\epsilon_{\mathrm{total}}$. We report the means of average match accuracy over three runs.
@@
$\epsilon_{\mathrm{token}}=1$ $0.37$ $\bm{0.40}$ $\bm{0.40}$ $0.45$ $\bm{0.49}$ $\bm{0.49}$ $0.45$ $0.49$ $0.49$ $0.45$ $0.49$ $0.49$
$\epsilon_{\mathrm{token}}=2$ $0.25$ $0.27$ $0.27$ $0.42$ $0.41$ $0.41$ $\bm{0.51}$ $0.50$ $0.49$ $\bm{0.51}$ $0.50$ $0.49$
$\epsilon_{\mathrm{token}}=5$ $0.12$ $0.12$ $0.12$ $0.25$ $0.27$ $0.27$ $0.38$ $0.38$ $0.38$ $0.46$ $0.46$ $0.44$
@@
$\epsilon_{\mathrm{token}}=1$ $0.28$ $0.37$ $0.45$ $0.28$ $0.38$ $0.46$ $0.28$ $0.38$ $0.46$ $0.28$ $0.38$ $0.46$
$\epsilon_{\mathrm{token}}=2$ $0.42$ $0.45$ $\bm{0.48}$ $0.48$ $0.55$ $\bm{0.57}$ $0.49$ $0.57$ $\bm{0.60}$ $0.49$ $0.57$ $0.60$
$\epsilon_{\mathrm{token}}=5$ $0.33$ $0.33$ $0.34$ $0.47$ $0.48$ $0.49$ $0.56$ $0.55$ $0.55$ $0.59$ $0.60$ $\bm{0.61}$
@@
Figure 3 shows the average match accuracy of baseline algorithms and our private algorithms under different total privacy guarantees ($\epsilon_{\mathrm{total}}$). Across different datasets and LLMs, we observe that DPSparseVoteRAG outperforms Non-RAG mostly under $\epsilon_{\mathrm{total}}\geq 10$ and approaches the upper bound of VoteRAG as we allow a larger privacy budget. This demonstrates that our algorithms enable us to exploit the external knowledge through RAG to improve the utility of QA tasks while ensuring a reasonable level of privacy.
@@
In Figure 3, we find that DPSparseVoteRAG consistently outperforms DPVoteRAG across different LLMs and datasets. DPSparseVoteRAG augments DPVoteRAG by utilizing the non-RAG LLM and the sparse vector technique so that it only spends a privacy budget for an output token requiring sensitive external knowledge. The consistently better performances of DPSparseVoteRAG suggest the importance of separately treating token generations for meaningful tokens, i.e., tokens requiring external knowledge, and for other general tokens in the privacy-constraint setting.
@@
We take a closer look at the effects of the hyperparameters in Table 1 with OPT (1.3B) on Trivia dataset under different total privacy budgets $\epsilon_{\mathrm{total}}$. We provide the detailed results, as in Table 1, with other LLMs in Appendix B.1.
Commonly between our private algorithms, we observe that the optimal $\epsilon_{\mathrm{token}}$ increases as we allow more total privacy budgets. Under a tight total privacy budget, large $\epsilon_{\mathrm{token}}$ allows our algorithms to only output a few meaningful tokens; thus, smaller $\epsilon_{\mathrm{token}}$ is preferable. Conversely, under a large total privacy budget, accurate token generation with large $\epsilon_{\mathrm{token}}$ is more important than having more tokens generated with small $\epsilon_{\mathrm{token}}$. Therefore, it is advised that we set $\epsilon_{\mathrm{token}}$ to be as large as possible to enable accurate token generations as long as it is small enough to allow the algorithms to generate a reasonably large number of tokens ($\approx 10$). Notice that DPSparseVoteRAG generally allows us to set larger $\epsilon_{\mathrm{token}}$ than DPVoteRAG under a fixed total privacy budget. This implies the benefit of DPSparseVoteRAG to save and spend a privacy budget cleverly—it can spend the saved privacy budget for generating important tokens for answering questions correctly.
With regard to the number of voters $m$, we generally see that more voters yield better utility with $\epsilon_{\mathrm{token}}=1$, but the number of voters has less effect on the utility with larger $\epsilon_{\mathrm{token}}$. This is due to the two distinct consequences of having more voters. More voters alleviate the effect of DP noise on the token histograms constructed in the algorithms. However, depending on the number of relevant documents to the question, there is a risk of having voters with irrelevant documents who can vote for the wrong tokens. The first consequence is more dominant particularly under small $\epsilon_{\mathrm{token}}$ while the second is more dominant under larger $\epsilon_{\mathrm{token}}$. Hence, $m$ should be set to balance these two consequences for achieving better per-token generation quality.